Determine a rule for the nth term of the sequence. Then find a17.
86, 82, 78, 74, 70 ... a17 =?

Answer by robertb(5830)

an = a1 + (n-1)d = 86 + (n-1)(-4) = 86 - 4n + 4 = 90 - 4n. Hence an = 90 - 4n.
Then a17 = 90 - 4*17 = 22.